OpenStudy (turbodog):

Find the differential of the function w=x^3sin(y^6z^3) This is what I got but its not correct. dw=3x^3y^6z^2cos(y^6z^3) dx=3x^2sin(y^6z^3) dy=6x^3y^5z^3cos(y^6z^3) Where am I going wrong?

OpenStudy (3mar):

First of all, this is a function that includes more than one variable, so when you are going to differentiate it, seclect one variable to take the derivative with respect to it.

OpenStudy (turbodog):

That is what I did, or was attempting to do anyway. In the case of x, I only took the derivative of x, treating the rest as constants

OpenStudy (3mar):

Good concept, but you wrote dw, although there is not w-variable in that equation!!

OpenStudy (turbodog):

dw is actually one of the questions I am supposed to answer, which has me confused.

OpenStudy (loser66):

They said \[dw= \dfrac{\partial w}{\partial x}+\dfrac{\partial w}{\partial y}+\dfrac{\partial w}{\partial z}\]

OpenStudy (loser66):

\(\dfrac{\partial w}{\partial x}= 3x^2sin(y^6z^3)\)

OpenStudy (loser66):

do the same with y and z.

OpenStudy (3mar):

Loser66 is right.

OpenStudy (turbodog):

Ah ok, thanks. that makes sense with respect to dw. However, the answer for x that you posted is the same I got but I was told it is incorrect

OpenStudy (loser66):

Not the same!!

OpenStudy (loser66):

Your answer: \(3x^3y^6z^2cos (y^6z^3)\) My answer: \(3x^2sin(y^6z^3)\)

OpenStudy (irishboy123):

shouldn't it be \(dw= \dfrac{\partial w}{\partial x} \color{red}{dx}+\dfrac{\partial w}{\partial y} \color{red}{dy}+\dfrac{\partial w}{\partial z} \color{red}{dz}\) ...which is called the **total differential**

OpenStudy (turbodog):

That was the answer I had for dw, but if you look at the attacked screen shot, my dx is \[3x ^{2}\sin(y ^{6}z ^{3}\]

OpenStudy (3mar):

hey @IrishBoy123 how do you change font colour?

OpenStudy (irishboy123):